How do I calculate this Equation?

Seventy-Two Inc., a developer of radiology equipment, has stock outstanding as follows: 80,400 shares of cumulative preferred 3% stock, $15 par, and 399,900 shares of $23 par common. During its first four years of operations, the following amounts were distributed as dividends: first year, $57,000 ; second year, $77,800 ; third year, $78,500 ; fourth year, $99,800 .

Calculate the dividends per share on each class of stock for each of the four years. (Preferred Stock & Common Stock are my two needed categories)
Answer #1

Annual preferred dividend = Number of preferred shares outstanding x Par value per preferred stock x Dividend rate

= 80,400 x 15 x 3%

= $36,180

Dividend Distribution Table
Year Paid to Preferred Stockholders Paid to Common Stockholders
1 36,180 57,000-36,180 = 20,820
2 36,180 77,800-36,180 = 41,620
3 36,180 78,500-36,180 = 42,320
4 36,180 99,800-36,180 = 63,620
Dividend Per Share Table
Year Preferred Stock Common Stock
1 36,180/80,400 = 0.45 20,820/399,900 = 0.05
2 36,180/80,400 = 0.45 41,620/399,900 = 0.10
3 36,180/80,400 = 0.45 42,320/399,900 = 0.11
4 36,180/80,400 = 0.45 63,620/399,900 = 0.16

Kindly comment if you need further assistance.

Thanks‼!